Tomato weights and Fertilizer (Raw Data, Software Required):
Carl the farmer has three fields of tomatoes, on one he used no fertilizer, in another he used organic fertilizer, and the third he used a chemical fertilizer. He wants to see if there is a difference in the mean weights of tomatoes from the different fields. The sample data for tomato-weights in grams is given below. Carl claims there is a difference in the mean weight for all tomatoes between the different fertilizing methods.

No Fertilizer Organic Fertilizer Chemical Fertilizer
123 112 115
119 127 141
95 138 143
97 133 134
94 140 129
120 114 134
114 126 135
118 134 129
129 123 113
129 143 148

The Test: Complete the steps in testing the claim that there is a difference in the mean weight for all tomatoes between the different fertilizing methods.

(a) What is the null hypothesis for this test?

H0:  μ1 = μ2 = μ3.

H0:  μ1μ2μ3.    

H0: At least one of the population means is different from the others.

H0:  μ3 > μ2 > μ1.


(b) What is the alternate hypothesis for this test?

H1:  μ1 = μ2 = μ3.

H1:  μ1μ2μ3.    

H1: At least one of the population means is different from the others.

H1:  μ3 > μ2 > μ1.


(c) Use software to get the P-value of the test statistic ( F ). Round to 4 decimal places unless your software automatically rounds to 3 decimal places.
P-value =

(d) What is the conclusion regarding the null hypothesis at the 0.05 significance level?

reject H0

fail to reject H0    


(e) Choose the appropriate concluding statement.

We have proven that all of the mean weights are the same.

There is sufficient evidence to conclude that the mean weights are different.    

There is not enough evidence to conclude that the mean weights are different.


(f) Does your conclusion change at the 0.10 significance level?

Yes

No    
Answer #1

Ans:

a)

H0:  μ1 = μ2 = μ3

b)

H1: At least one of the population means is different from the others.

c)p-value=0.0040

d)As,p-value<0.05,so reject H0

e)There is sufficient evidence to conclude that the mean weights are different.    

f)No,as p-value<0.10,so reject H0

ANOVA :

Group 1 Group 2 Group 3 Total
Sum 1138 1290 1321 3749
Count 10 10 10 30
Mean, Sum/n 113.8 129 132.1
Sum of square, Ʃ(xᵢ-x̅)² 1657.6 1002 1142.9
Standard deviation 13.5712 10.5515 11.2689
Number of treatment, k = 3
Total sample Size, N = 30
df(between) = k-1 = 2
df(within) = N-k = 27
df(total) = N-1 = 29
SS(between) = (Sum1)²/n1 + (Sum2)²/n2 + (Sum3)²/n3 - (Grand Sum)²/ N = 1918.4667
SS(within) = SS1 + SS2 + SS3 = 3802.5000
SS(total) = SS(between) + SS(within) = 5720.9667
MS(between) = SS(between)/df(between) = 959.2333
MS(within) = SS(within)/df(within) = 140.8333
F = MS(between)/MS(within) = 6.8111
p-value = F.DIST.RT(6.8111, 2, 27) = 0.0040
Source of Variation SS df MS F P-value
Between Groups 1918.4667 2 959.2333 6.8111 0.0040
Within Groups 3802.5000 27 140.8333
Total 5720.9667 29